OpenStudy (anonymous):

. Select the set of equations that represents the following situation: Mary invested one amount at 7% simple interest, and a second amount at 5% interest, earning $29.80 in one year. If she had switched the amounts, she would have earned $35.00. What were the two amounts?

OpenStudy (campbell_st):

let x = amount 1 and y = amount 2 using I = Prn then the equations are..... 7% = 0.07 5% = 0.05 n = 1 0.07x + 0.05y = 29.80 (1) 0.05x + 0 07y = 35.00 (2) solve by elimination multiply equation (1) by 5 and subtract equation (2) multiplied by 7 0.35x + 0.25y = 149 - 0.35x + 0.49y = 245 ------------------------ -0.24y = -96 y = $400 substitute into equation (1) to find x 0.07x + 0.05(400) = 29.80 0.07x = 9.80 x = $140 the investments where y = $400 and x = $140